OpenStudy (anonymous):

please explain how you got the answer:) medals Add. –32.7 + (–0.2) A. –32.9 B. –32.5 C. 32.5 D. 32.9

OpenStudy (cggurumanjunath):

-32.7+(-0.2)=-32.7-0.2=-(32.7+0.2)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Take -32.7 Add - 0.2 This also looks like -32.7 - 0.2 Subtract in the tenths place -32.9 Does that make sense?

OpenStudy (paki):

–32.7 + (–0.2) = -32.7-0.2= -32.5

OpenStudy (anonymous):